A mixture of 30 pounds of candy sells for $1.10 a pound. The mixture consists of chocolates worth $1.50 a pound and chocolates worth 90¢ a pound. How many pounds of the $1.50 chocolate were used to make the mixture?
If x represents the pounds of $1.50 chocolates used, then which of the following represents the pounds of 90¢ chocolates used?

Answer by mananth(16949)

price ---------------- quantity
candy I 1.5----------------x lb
candy II0.9---------------- 30-xlb
Total 1.10% ----------------30
1.5x+0.9(30-x)= 1.1*30
1.5x+27 -0.9x=33
1.5x-0.9x=33-27
0.6x= 6
/0.6
x=10lb $1.50
20lb $0.90